PLEASE TELL ME HOW I CAN SOLVE IT,I AM STUCK
(IF 16*A^2*64=(4*B)^2 AND A AND B ARE POSITIVE INTEGERS,THEN B IS HOW MANY TIMES GREATER THAN A?
PLEASE EXPLAIN ME WHAT ARE THEY ASKING FOR AND HOW CAN I SOLVE IT?

Found 2 solutions by ptaylor, MathTherapy:
Answer by ptaylor(2198)

16*A^2*64=(4*B)^2
16*(A^2)*64=16*B^2 Divide both sides by 16
64*A^2=B^2 Take the square root of both sides
8A=B
The problem asks to determine how many times greater B is than A.
If it takes 8 A's to make 1 B, the B must be 8 times greater than A
Hope this helps--ptaylor

Answer by MathTherapy(10552)


PLEASE TELL ME HOW I CAN SOLVE IT,I AM STUCK
(IF 16*A^2*64=(4*B)^2 AND A AND B ARE POSITIVE INTEGERS,THEN B IS HOW MANY TIMES GREATER THAN A?
PLEASE EXPLAIN ME WHAT ARE THEY ASKING FOR AND HOW CAN I SOLVE IT?

B is 7 times GREATER THAN A
===================
If you need a complete and detailed solution, let me know!!
Send comments, “thank-yous,” and inquiries to “D” at MathMadEzy@aol.com.
Further help is available, online or in-person, for a fee, obviously.
For FREE info and answers to questions about the ASVAB exam, the NYS 3 – 8 city/state wide exams,GENERAL
MATH and HOMEWORK QUESTIONS, or MATH QUESTIONS related to the Regents Integrated Algebra,
Regents Geometry, Regents Algebra 2/Trigonometry, SHSAT, COOP/HSPT/TACHS, PSAT, SAT, ACT, SSAT/ISEE,
GRE, CLEP, and the TASC/GED, you can visit: http://asvabstudyzone.freeforums.net/.